# Towards Generalizable AI-Generated Image Detection via Image-Adaptive Prompt Learning

### v2

@@ 4.3 Main Results
We report the accuracy and average precision on the UniversalFakeDetect dataset in Tab. 1 and Tab. 2, respectively. Our proposed method achieves SOTA results with 95.61% mAcc and 99.32% mAP across 19 diverse sub-test sets, outperforming all existing methods. In terms of mAcc, we surpass the SOTA methods C2P-CLIP by 1.82%. In terms of mAP, we surpass the SOTA methods RINE by 0.53%. Furthermore, compared with the baseline UniFD, our method brings substantial improvements of 14.23% in mAcc and 9.18% in mAP. These gains clearly demonstrate that our model, when trained solely on GAN-generated data [18], can generalize robustly to a wide range of challenging testing scenarios. Beyond the overall metrics, our method ranks first or second on 12 out of 19 sub-test sets in terms of mAP. While previous approaches may excel on specific generative models but degrade considerably on others, our method maintains consistently strong performance across almost all synthesis types, highlighting its superior generalization ability. This is mainly attributed to our adaptive design, which enables the model to capture diverse artifacts with minimal bias toward any specific generative family.
@@
Table 4: Effectiveness of each proposed component of our method.
Component mAcc mAP
Baseline 84.19 93.52
+MLP-based Adapters 89.02 97.66
+Learnable Tokens 91.83 98.04
+Image-Adaptive Prompts 94.88 98.72
+Optimal View Selection 95.61 99.32
Table 5: Ablation of Image-Adaptive Prompt Learning. RAcc and FAcc denote mAcc for real and fake images, respectively. Best results are highlighted in blue.
FS cond. IS cond. Gated Mech. TTT. Conf. Sel. mAcc mAP RAcc FAcc
@@
Table 7: Ablation of the number of learnable tokens $N_{t}$.
@@
For the GenImage dataset, we report the overall accuracy results in Tab. 3. Our method achieves 96.7% accuracy on the full test set, significantly surpassing recent methods CO-SPY and AIDE by 12.0% and 9.8%, respectively. In comparison with SOTA approaches such as C2P-CLIP and SAFE, our method further improves the mean mAcc by 0.9% and 1.1%. Moreover, it outperforms the baseline UniFD by a large margin of 17.2% in mAcc. These results demonstrate that our method, trained on a diffusion-based synthetic dataset [51], exhibits strong generalization capability across diverse types of testing manipulations. Beyond the overall metrics, our method ranks top-2 on 7 out of 8 sub-test sets, highlighting its stability and generalization across diverse scenarios. Results on both UniversalFake and GenImage datasets confirm the consistent performance of our method across various forgeries.
@@ 4.4 Ablation Study
As shown in Tab. 5, we first verify the effectiveness of the four main modules of our method. It can be observed that the MLP-based adapter, learnable tokens, image-adaptive prompt, and Optimal View Selection all significantly improve the detection performance. Using them together achieves the best performance. Importantly, even with only simple MLP-based adapters and learnable tokens, our method surpasses prior approaches with heavy architectural changes, such as Fatformer. As shown in Tab. 5, we progressively evaluate the contributions of forgery-specific condition (FS cond.), image-specific condition (IS cond.), the gated mechanism (Gated Mech.), Test-Time Token Tuning (TTT), and Confidence Selection (Conf. Sel.). Each component consistently improves performance, with the full configuration achieving the highest accuracy, average precision, real accuracy, and fake accuracy. As shown in Tab. LABEL:tab:6, we explore several strategies for inserting the MLP-based adapter, including placing it in the first 6 layers, the last 6 layers, and evenly distributing it. The results show that evenly distributed insertion achieves the best performance. Moreover, inserting the adapter into 6 layers already yields satisfactory performance, and further increasing the number of inserted layers brings only marginal improvements. As shown in Tab. LABEL:tab:7, we analyze the effect of varying the number of learnable tokens ($N_{t}$) in our method. Under the current experimental setup, the image-adaptive prompt in the first layer is replaced with learnable tokens. The results show that using 9 tokens achieves the best performance, indicating a balanced trade-off between model capacity and efficiency. Performance decreases slightly when using more or fewer learnable tokens. As shown in Tab. LABEL:tab:8, we conduct experiments on the choice of loss used in Test-Time Token Tuning. It can be observed that using averaged entropy losses yields better results. This may be because averaged entropy loss could be regarded as the sum of pointwise entropy loss and the KL divergence term [48]. The KL divergence term encourages consistent predictions among selected views. Improving prediction consistency across different views in forgery detection can reduce the interference of local misjudgments and enhance the model’s ability to detect images with ambiguous boundaries or low levels of forgery. As shown in Tab. LABEL:tab:9, we present the effect of varying the number of tuning steps ($T$) on model performance. As $T$ increases from 1 to 3, both accuracy Acc and AP steadily improve. Finally, we set $T=2$ for a balanced trade-off between efficiency and effectiveness.
As shown in Fig. 5, we visualize the Grad-CAM [47] of without and with Image-Adaptive Prompt Learning (IAPL). It could be observed that IAPL enables the model to focus more precisely on forgery-related regions, whereas the baseline tends to attend to irrelevant areas.
@@ 5 Conclusion
In this paper, we propose Image-Adaptive Prompt Learning (IAPL), a novel strategy for generalizable AI-generated image detection. Unlike prior fine-tuning approaches that rely on fixed parameters and struggle with unseen forgeries, IAPL dynamically adapts prompts to each input image via a Conditional Information Learner and Test-Time Token Tuning. The former captures both forgery-specific and image-specific cues, while the latter optimizes adaptive tokens during inference by enforcing consistency across augmented views, ensuring alignment with the characteristics of each image. Extensive experiments demonstrate that IAPL achieves superior robustness and SOTA performance.
Limitation. Although our method achieves satisfactory mAcc on both the UniversalFakeDetect and GenImage datasets, its performance on certain generation methods still leaves room for improvement.

### v3

@@ 4.3 Main Results
We report the accuracy and average precision on the UniversalFakeDetect dataset in Tab. 1 and Tab. 2, respectively. Our proposed method achieves SOTA results with 95.61% mAcc and 99.32% mAP across 19 diverse sub-test sets, outperforming all existing methods. In terms of mAcc, we surpass the SOTA methods C2P-CLIP by 1.82%. In terms of mAP, we surpass the SOTA methods RINE by 0.53%. Furthermore, compared with the baseline UniFD, our method brings substantial improvements of 14.23% in mAcc and 9.18% in mAP. These gains clearly demonstrate that our model, when trained solely on GAN-generated data [goodfellow2020generative], can generalize robustly to a wide range of challenging testing scenarios. Beyond the overall metrics, our method ranks first or second on 9 out of 19 sub-test sets in terms of mAcc. While previous approaches may excel on specific generative models but degrade considerably on others, our method maintains consistently strong performance across almost all synthesis types, highlighting its superior generalization ability. This is mainly attributed to our adaptive design, which enables the model to capture diverse artifacts with minimal bias toward any specific generative family.
@@
For the GenImage dataset, we report the overall accuracy results in Tab. 3. Our method achieves 96.7% accuracy on the full test set, significantly surpassing recent methods ATTSD and MiraGe by 6.0% and 4.1%, respectively. In comparison with the SOTA approach C2P-CLIP, our method further improves the mean mAcc by 0.9%. Moreover, it outperforms the baseline UniFD by a large margin of 17.2% in mAcc. These results demonstrate that our method, trained on a diffusion-based synthetic dataset [song2020score], exhibits strong generalization capability across diverse types of testing manipulations. Beyond the overall metrics, our method ranks first or second on 7 out of 8 sub-test sets, highlighting its stability and generalization across diverse scenarios. Results on both UniversalFake and GenImage datasets confirm the consistent performance of our method across various forgeries and generators.
Table 4: Effectiveness of each proposed component. RAcc and FAcc denote mAcc for real and fake images.
Component mAcc mAP RAcc FAcc
Baseline 84.19 93.52 87.24 81.15
+MLP-based Adapters 89.02 97.66 86.76 91.28
+Learnable Tokens 91.83 98.04 96.79 86.87
+Image-Adaptive Prompts 94.88 98.72 97.07 92.68
+Optimal Input Selection 95.61 99.32 96.37 94.84
Table 5: Ablation of Image-Adaptive Prompt Learning. Default configurations are indicated with a background in blue.
FSC. GEC. LSF. TTT. CFS. mAcc mAP RAcc FAcc
@@
Table 7: Number of layers $N_{t}$ with learnable tokens.
@@ 4.4 Ablation Study
As shown in Tab. 4, we first verify the effectiveness of the four main modules of our method. It can be observed that the MLP-based adapter, learnable tokens, image-adaptive prompt, and Optimal Input Selection all significantly improve the detection performance. Using them together achieves the best performance. Importantly, even with only simple MLP-based adapters and learnable tokens, our method surpasses prior approaches with heavy architectural changes, such as Fatformer. As shown in Tab. 5, we progressively evaluate the contributions of forgery-specific condition (FSC.), general condition (GEC.), the learnable scaling factor (LSF.), Test-Time Token Tuning (TTT.), and Confidence Selection (CFS.). Each component consistently improves performance, with the full configuration achieving the highest accuracy, average precision, real accuracy, and fake accuracy. As shown in Tab. LABEL:tab:6, the three numbers in the config represent the starting layer, the ending layer, and the interval between inserted adapters. The results indicate that evenly inserting adapters into six layers yields better performance than placing them in the first six layers, the last six layers, or inserting them into more layers. Moreover, inserting the adapter into 6 layers already yields satisfactory performance, and further increasing the number of inserted layers brings only marginal improvements. As shown in Tab. LABEL:tab:7, we validated the number of encoder blocks $N_{t}$ where learnable tokens are employed. Under the current experimental setup, the image-adaptive prompt in the first layer is replaced with learnable tokens. The results show that using 9 tokens achieves the best performance, indicating a balanced trade-off between model capacity and efficiency. Performance decreases slightly when using more or fewer learnable tokens. As shown in Tab. LABEL:tab:8, we conduct experiments on the choice of loss used in Test-Time Token Tuning. It can be observed that using averaged entropy losses yields better results. This may be because the averaged entropy loss could be regarded as the sum of pointwise entropy loss and the KL divergence term [sheng2025r]. The KL divergence term encourages consistent predictions among selected views. Improving prediction consistency across different views in forgery detection can reduce the interference of local misjudgments and enhance the model’s ability to detect images with ambiguous boundaries or low levels of forgery. As shown in Tab. LABEL:tab:9, we present the effect of varying the number of tuning steps ($T$) on model performance. As $T$ increases from 1 to 3, both accuracy Acc and AP steadily improve. Finally, we set $T=2$ for a balanced trade-off between efficiency and effectiveness. As shown in Fig. 5, we visualize the Grad-CAM [selvaraju2017grad] without and with IAPL. It could be observed that IAPL enables the model to focus more precisely on forgery-related regions, whereas the baseline tends to attend to irrelevant areas. Refer to the supplementary for more visualizations and experiments.
@@ 5 Conclusion
In this paper, we propose Image-Adaptive Prompt Learning (IAPL), a novel strategy for generalizable AI-generated image detection. Unlike prior fine-tuning approaches that rely on fixed parameters and struggle with unseen forgeries, IAPL dynamically adapts prompts to each input image via a Conditional Information Learner and Test-Time Token Tuning. The former captures both forgery-specific and general cues, while the latter optimizes adaptive tokens during inference by enforcing consistency across augmented views, ensuring alignment with the characteristics of each image. Extensive experiments demonstrate that IAPL achieves superior robustness and SOTA performance.